OpenStudy (anonymous):

find the directional derivative of f at point p f(x,y)= arctan (y/x) p(4,-4) u=2i-3j

OpenStudy (kinggeorge):

First step, find the unit vector in the direction of u.

OpenStudy (anonymous):

but isn't the unit vector already given to you?

OpenStudy (anonymous):

or is it just 1/||u||, in which case its 1/4 (2i-3j) which would be 1/2i - 3/4j?

OpenStudy (anonymous):

wait.. 4 + 9 is not 16...

OpenStudy (kinggeorge):

So let \(\vec u_0\) be directional derivative in the direction of \(\vec u\). Then \[\vec u_0={2 \over \sqrt{13}}\vec i - {3 \over \sqrt{13}}\vec j\]

OpenStudy (anonymous):

2/sqrt(13)i+3/sqrt(13)j??

OpenStudy (anonymous):

i'm not quite sure how to find Fx or Fy and btw, HI GEORGE haha

OpenStudy (kinggeorge):

It should be -3 since your original vector has a -3.

OpenStudy (anonymous):

ohh that's right..

OpenStudy (kinggeorge):

Hi :) I had to go to bed last night so I missed the question you tagged me in.

OpenStudy (anonymous):

no worries haha

OpenStudy (kinggeorge):

Anyways, next step is to find \(f_x\) and \(f_y\).

OpenStudy (anonymous):

this is my guess: Fx = -(sec^2)^-2(xy^-1)y^-1 Fy= -(sec^2)^-2(xy^-1)x

OpenStudy (kinggeorge):

We can use the fact that \[{ d \over dx}\;\;\arctan(x)={1 \over {1+x^2}}\]

OpenStudy (anonymous):

wait, is that a rule?

OpenStudy (kinggeorge):

Thus, \[\large f_x={\partial \over \partial x} \;\;\arctan\left({y \over x}\right)={1 \over 1+\left({y \over x}\right)^2}\cdot ({\partial \over \partial x}\;\;\; {y \over x})={-y \over {x^2+y^2}}\]

OpenStudy (kinggeorge):

And yes that is a rule. Feel free to use it whenever.

OpenStudy (kinggeorge):

Likewise, \[\large f_y= {\partial \over \partial y}\;\;\arctan\left({y \over x}\right)={1 \over 1+\left({y \over x}\right)^2}\cdot \left({\partial \over \partial y} \;\;\;{y \over x}\right)={x \over x^2+y^2}\](I skipped some steps in the simplification)

OpenStudy (anonymous):

are you multiplying it by ∂/∂y yx because of chain rule?

OpenStudy (kinggeorge):

Yes. In this case, we don't have to do any complicated sort of chain rule like we did last night.

OpenStudy (anonymous):

so then gradf = -4/32 x (2/sqrt(13) + 4/32 x (-2/sqrt(13)?!

OpenStudy (kinggeorge):

In the second term, you should have a -3 instead of -2. However, that is the directional derivative. The gradient is slightly different.

OpenStudy (anonymous):

oops, that was a typo, but aren't i supposed to get the directional derivative by getting the gradient and dot product'ing it with the unit vector?

OpenStudy (kinggeorge):

That's basically what we did here, just in a more direct way. We calculated the coefficients of the unit vector first, and then the coefficients of the gradient. By multiplying them together the way we did, it was like we found the gradient and used the dot product on the unit vector.

OpenStudy (anonymous):

ohh i see what you're saying, so you weren't correcting my answer, just my term okay that makes sense

OpenStudy (kinggeorge):

Precisely.

OpenStudy (anonymous):

ohh okay, cool! thanks georige!!

OpenStudy (kinggeorge):

You're welcome.